OpenStudy (anonymous):

Combinations: HELP! TELL ME WHAT IM DOING WRONG:( A schools academic team consists of 7 students, but only two can participate in this years Knowledge slame. How many different 2- person teams can be formed from 7 students? 7! -- (7-2)!2! next step 7! -- (5)!2! so 7 x 6 x 5 x 4 x 3 x 2 ----------------- 5 x 4 x 3 x 2 x 2 ??? I don't know where to go from here, my answer is always wrong no matter what I do from here. I followed the combinations formal but I don't know! Help :(

OpenStudy (anonymous):

Never mind, I just got the answer! :) 21!!

OpenStudy (anonymous):

\[\frac{ 7! }{ 5!2! }\] \[\frac{ 7*6*5*4*3*2*1 }{ 2*1*5*4*3*2*1 }\] \[\frac{ 7*6 }{ 2*1 }\] \[\frac{ 42 }{ 2 }\]

OpenStudy (anonymous):

\[21\]

OpenStudy (amistre64):

\[nCr=\frac{nPr}{r!}\]that tends to help me simplify those things in my own mind